# TrajLearn: Trajectory Prediction Learning using Deep Generative Models

### v1

@@ 6.1. Motivation
  • Efficient Resource Utilization: By allocating larger hexagons to low-activity areas, the approach reduces unnecessary computational and memory overhead, focusing resources on regions where fine-grained predictions are most beneficial.
@@
  • Improved Data Representation: By matching the map’s resolution to the inherent data distribution, hierarchical maps provide a more accurate and meaningful representation of spatial dynamics, enhancing overall model performance.
@@ 6.2. Approach
2 $H\leftarrow\texttt{Tessellate}(M,R_{\text{min}})$;
@@
5       if termination_condition_fn($H$) then
@@
[…]
@@ 6.4. Experimental Results
Figure 16 illustrates the distinction between fixed-resolution and hierarchical tessellations for the Ho-GeoLife dataset. The hierarchical map achieves greater spatial specificity in high-density areas, facilitating more accurate and localized trajectory predictions without incurring significant computational overhead in less active regions.
@@ 7.1.1. Computer Vision
Trajectory prediction in computer vision involves predicting the future movement of objects in a scene over time. They rely on camera-generated video frames, where trajectories can be represented by $(x,y)$ coordinates within the frame (Huang et al., 2019; Cong et al., 2022; Li et al., 2022). The focus is on trajectory prediction for autonomous driving (Ren et al., 2021; Song et al., 2020), pedestrian mobility prediction at a small scale (Sun et al., 2022; Duan et al., 2022), or predicting human–human and human–vehicle interactions (Huang et al., 2019; de Brito et al., 2021). Despite their effectiveness, computer vision-based approaches often encounter significant challenges that limit their scalability and applicability to real-world scenarios. Primarily, the reliance on camera systems imposes constraints such as limited fields of view, which restrict the ability to capture comprehensive spatial dynamics in expansive environments. Additionally, these methods frequently depend on visual features like optical flow and bounding box detections, which are not available in datasets lacking visual information. In contrast, our approach circumvents these limitations by only deriving its input from GPS data.
@@ 7.1.2. Mobile Data Analysis
Microscopic Analysis. Microscopic analysis focuses on individual-level mobility prediction and is more closely related to our work. First attempts to address the problem considered statistical methods, such as matrix factorization (Cheng et al., 2012; Li et al., 2015; Lian et al., 2014) and Markov chain (Cheng et al., 2013; Mathew et al., 2012; Shi et al., 2019). However, these approaches often struggle to capture human mobility’s complex sequential and periodic features in trajectories. Deep Learning advances yield specialized models for sequential trajectory modeling. Particularly, methods based on RNNs have demonstrated good performance. For instance, Liu et al. (Liu et al., 2016) proposed Spatial Temporal Recurrent Neural Networks (ST-RNN), a method that extends RNN to model temporal and spatial contexts. Unlike traditional methods such as Markov Chains, Factorization Models, and standard RNNs, which struggle with continuous time intervals, geographical distances, or sparse data, ST-RNN employs time-specific and distance-specific transition matrices to capture dynamic temporal and spatial dependencies. By leveraging a linear interpolation method, it mitigates data sparsity issues and enhances predictive accuracy. Feng et al. (Feng et al., 2018) presented DeepMove, a attentional recurrent networks designed to predict human mobility from sparse and lengthy trajectory data. It addresses challenges like complex sequential dependencies and multi-level periodicity in movement patterns by combining a multi-modal embedding RNN for feature representation with a historical attention module to leverage relevant historical patterns. Experiments on three real-world datasets demonstrate that DeepMove outperforms traditional models by over 10 % in accuracy, while its attention mechanism offers interpretability by highlighting key historical influences. While these models perform well, they struggle to handle sparse and inaccurate trajectory data.
@@
Lian et al. (Lian et al., 2020) introduce GeoSAN (Geography-Aware Sequential Recommendation based on Self-Attention Network), a model designed to enhance sequential location recommendation by effectively incorporating geographical information and addressing data sparsity. GeoSAN employs a self-attention-based geography encoder to embed spatial proximity and clustering phenomena, alongside a novel loss function that uses importance sampling to prioritize informative negative samples. Additionally, geography-aware negative samplers are introduced to enhance the informativeness of training data. Also, Yang et al. (Yang et al., 2020) (Flashback), Luo et al. (Luo et al., 2021) (STAN), and Xue, Hao, et al. (Xue et al., 2021)(MobTCast) introduced models crafted for handling sparse user mobility data. However, these models are designed for the Next POI prediction problem, which, as explained in section 1, is a different problem. Earlier studies investigated trajectory prediction using general GPS logs rather than POI check-ins, yet mainly concentrated on specific instances of the problem. For example, Jiang et al. (Jiang et al., 2018) employed a seq2seq model to predict very short-term human trajectories triggered by big rare events or disasters. Sadri et al. (Sadri et al., 2018) presented a model for predicting a user’s afternoon trajectory, given their morning trajectory. Amichi et al. (Amichi et al., 2021) proposed to first predict the purpose of visiting a location and, given that, to predict the next location where the individual will be. The paper (Yan et al., 2023) introduces PreCLN, a pre-trained contrastive learning framework for vehicle trajectory prediction that leverages a dual-view approach combining hierarchical map gridding and road network mapping to capture spatiotemporal dependencies. It employs a Transformer-based trajectory encoder to model long-term relationships and integrates three auxiliary pre-training tasks—trajectory imputation, destination prediction, and trajectory-user linking—to enhance representation learning. Experimental results on large-scale trajectory datasets demonstrate significant improvements over state-of-the-art baselines showcasing its ability to handle complex trajectory data and improve prediction accuracy for smart transportation applications. However, this approach requires time, user embeddings, and three auxiliary pre-trained tasks to train the model, increasing information demands and complexity. These models are unsuitable for addressing the trajectory prediction problem.
@@ 8. Ethical Implications
Using deep generative models for trajectory prediction presents several ethical considerations that should be carefully addressed.
Privacy and Data Protection. Using trajectory data raises concerns about individual privacy and the potential for re-identification of individuals. To protect the privacy of individuals all datasets used in the experimental evaluation have been anonymized and aggregated. The original datasets are publicly available and are free of use (for research intent purposes), as outlined by their respective curators, and were deemed to be collected with proper informed consent and ethical approvals. Moreover, they have undergone the necessary preprocessing to meet specific research requirements. We also followed all terms and conditions of use as specified by the dataset providers, including properly attributing their work.
Other Ethical Consideration. Deep generative models pose challenges, such as the generation of plausible but incorrect data in certain scenarios. In our process of designing and developing our model, we have made a conscientious effort to proactively address potential ethical considerations. We have been diligent in adopting responsible practices to the best of our knowledge and capacity. By doing so, we aim to ensure that deploying deep generative models for trajectory prediction aligns with societal values and prioritizes the well-being of individuals and communities.
@@ 9. Conclusions
We focused on the problem of trajectory prediction and proposed TrajLearn, a trajectory deep generative model that has shown remarkable results in predicting the future path of a trajectory across various real-world datasets. Our model was trained from scratch on large-scale, higher-order mobility flow datasets that represent trajectories as sequences of hexagons on a hex-tessellated map. By utilizing higher-order mobility flows notable data simplification, while preserving essential spatial and temporal information, and incorporating a constrained beam search strategy, TrajLearn achieves significant advancements over state-of-the-art methods, with performance improvements of up to 40%. Additionally, our development of mixed-resolution maps demonstrates the model’s adaptability and practicality for diverse applications, from urban planning to autonomous navigation. Our extensive empirical evaluations across three real-world datasets, Ho-Porto, Ho-Rome, and Ho-GeoLife, demonstrate that TrajLearn consistently outperforms strong baselines, with Accuracy and BLEU score improvements of up to 40%. These results not only validate the robustness of our model but also highlight its competitive performance across different resolutions, showcasing its versatility. We believe our proposed approach and model for trajectory prediction can find many valuable applications and have a broad impact.

### v2

@@ 5.7.1. Hexagon Representation and Mapping to GPS Points
[…]
Diagram illustrating the conversion process from hexagon-based trajectory predictions to GPS coordinates. The figure shows a tessellated map of hexagons, where each hexagon is mapped to its centroid (GPS point).
@@ 5.7.2. Experimental Visualization
[…]
A map visualization depicting a trajectory from the GeoLife dataset. The image overlays several elements on a map: input GPS points outlining the trajectory, ground truth historical trajectory points, markers showing the centroids of hexagons predicted by TrajLearn, and a final GPS prediction obtained by map-matching the hexagon centroids. This illustrates the transformation from raw trajectory data to hexagon-based predictions and back to GPS coordinates.
@@ 6.1. Motivation
[…]
A multi-panel diagram illustrating the construction of a hierarchical map. Panel (a) displays an initial tessellation where a blue hexagon indicates a high-activity area. Panel (b) shows the subdivision of the blue hexagon into seven smaller hexagons for enhanced spatial precision. Panel (c) identifies an additional high-activity area with a red hexagon. Panel (d) presents the final tessellation with varying resolutions across the map, featuring smaller hexagons in regions of high activity.
@@
[…]
An illustrative diagram depicting the creation of a mix-resolution map. In panel (a), a tessellated map is shown with two red hexagons marking busy areas. Panel (b) demonstrates that these busy hexagons are subdivided into smaller hexagons. Panel (c) presents the final tessellation, where different areas of the map have varying hexagon sizes corresponding to different resolutions.
@@
  • Efficient Resource Utilization: By allocating larger hexagons to low-activity areas, the approach reduces computational and memory overhead, focusing resources on regions where fine-grained predictions are most beneficial.
  • Improved Data Representation: By matching the map’s resolution to the inherent data distribution, hierarchical maps provide a more accurate and meaningful representation of spatial dynamics, enhancing overall performance.
@@ 6.2. Approach
2 $H\leftarrow\mathtt{Tessellate}(M,R_{\text{min}})$;
@@
5       if $\mathtt{termination\_condition\_fn}$($H$) then
@@
9             if $\mathtt{splitting\_condition\_fn}$($h$) and $\mathtt{Resolution}$($h$) $<$ $R_{\text{max}}$ then
@@ 6.4. Experimental Results
[…]
A two-panel comparison of tessellated maps for the Ho-GeoLife dataset. Panel (a) displays a fixed resolution tessellation at resolution 7, showing movement density across the map. Panel (b) presents a hierarchical tessellation using resolutions 7, 8, and 9, where high-activity areas are depicted with finer granularity to enable more precise trajectory predictions.
Figure 16(b) illustrates the distinction between fixed-resolution and hierarchical tessellations for the Ho-GeoLife dataset. The hierarchical map achieves greater spatial specificity in high-density areas, facilitating more accurate and localized trajectory predictions without incurring significant computational overhead in less active regions.
@@ 7.1.1. Computer Vision
Trajectory prediction in computer vision involves predicting the future movement of objects in a scene over time. They rely on camera-generated video frames, where trajectories can be represented by $(x,y)$ coordinates within the frame (Cong et al., 2022; Li et al., 2022). The focus is on trajectory prediction for autonomous driving (Ren et al., 2021; Song et al., 2020), pedestrian mobility prediction at a small scale (Sun et al., 2022; Duan et al., 2022; Cao et al., 2021), or predicting human–human and human–vehicle interactions (Huang et al., 2019; de Brito et al., 2021). Despite their effectiveness, computer vision-based approaches often encounter significant challenges that limit their scalability and applicability to real-world scenarios. Primarily, the reliance on camera systems imposes constraints such as limited fields of view, which restrict the ability to capture comprehensive spatial dynamics in expansive environments. Additionally, these methods frequently depend on visual features like optical flow and bounding box detections, which are not available in datasets lacking visual information. In contrast, our approach circumvents these limitations by only deriving its input from GPS data.
@@ 7.1.2. Mobile Data Analysis
Microscopic Analysis. Microscopic analysis focuses on individual-level mobility prediction and is more closely related to our work. First attempts to address the problem considered statistical methods, such as matrix factorization (Cheng et al., 2012; Li et al., 2015; Lian et al., 2014) and Markov chain (Cheng et al., 2013; Mathew et al., 2012; Shi et al., 2019). However, these approaches often struggle to capture human mobility’s complex sequential and periodic features in trajectories. Deep Learning advances yield specialized models for sequential trajectory modeling. Particularly, methods based on RNNs have demonstrated good performance. For instance, Liu et al. (Liu et al., 2016) proposed Spatial Temporal Recurrent Neural Networks (ST-RNN), a method that extends RNN to model temporal and spatial contexts. Unlike traditional methods such as Markov Chains, Factorization Models, and standard RNNs, which struggle with continuous time intervals, geographical distances, or sparse data, ST-RNN employs time-specific and distance-specific transition matrices to capture dynamic temporal and spatial dependencies. By leveraging a linear interpolation method, it mitigates data sparsity issues and enhances predictive accuracy. Feng et al. (Feng et al., 2018) presented DeepMove, an attentional recurrent network designed to predict human mobility from sparse and lengthy trajectory data. It addresses challenges like complex sequential dependencies and multi-level periodicity in movement patterns by combining a multi-modal embedding RNN for feature representation with a historical attention module to leverage relevant historical patterns. Experiments on three real-world datasets demonstrate that DeepMove outperforms traditional models by over 10 % in accuracy, while its attention mechanism offers interpretability by highlighting key historical influences. While these models perform well, they struggle to handle sparse and inaccurate trajectory data.
@@
Lian et al. (Lian et al., 2020) introduce GeoSAN (Geography-Aware Sequential Recommendation based on Self-Attention Network), a model designed to enhance sequential location recommendation by effectively incorporating geographical information and addressing data sparsity. GeoSAN employs a self-attention-based geography encoder to embed spatial proximity and clustering phenomena, alongside a novel loss function that uses importance sampling to prioritize informative negative samples. Additionally, geography-aware negative samplers are introduced to enhance the informativeness of training data. Also, Yang et al. (Yang et al., 2020) (Flashback), Luo et al. (Luo et al., 2021) (STAN), and Xue, Hao, et al. (Xue et al., 2021)(MobTCast) introduced models crafted for handling sparse user mobility data. However, these models are designed for the Next POI prediction problem, which, as explained in section 1, is a different problem. Earlier studies investigated trajectory prediction using general GPS logs rather than POI check-ins, yet mainly concentrated on specific instances of the problem. For example, Jiang et al. (Jiang et al., 2018) employed a seq2seq model to predict very short-term human trajectories triggered by big rare events or disasters. Sadri et al. (Sadri et al., 2018) presented a model for predicting a user’s afternoon trajectory, given their morning trajectory. Amichi et al. (Amichi et al., 2021) proposed to first predict the purpose of visiting a location and, given that, to predict the next location where the individual will be. The paper (Yan et al., 2023) introduces PreCLN, a pre-trained contrastive learning framework for vehicle trajectory prediction that leverages a dual-view approach combining hierarchical map gridding and road network mapping to capture spatial-temporal dependencies. It employs a Transformer-based trajectory encoder to model long-term relationships and integrates three auxiliary pre-training tasks—trajectory imputation, destination prediction, and trajectory-user linking—to enhance representation learning. Experimental results on large-scale trajectory datasets demonstrate significant improvements over state-of-the-art baselines showcasing its ability to handle complex trajectory data and improve prediction accuracy for smart transportation applications. However, this approach requires time, user embeddings, and three auxiliary pre-trained tasks to train the model, increasing information demands and complexity. These models are unsuitable for addressing the trajectory prediction problem.
@@ 7.2.1. Transformers
Limitations of Conventional Generative Models for Trajectory Prediction. While deep generative models have achieved significant success in various applications, traditional time series forecasting methods, even those leveraging transformer architectures, are not directly applicable to our trajectory prediction problem. This is primarily because conventional methods typically predict continuous scalar or vector values at fixed intervals (Miller et al., 2024; Benidis et al., 2022), whereas our approach abstracts raw trajectory data into sequences of discrete spatial units via map tessellation, effectively mitigating issues of GPS noise and data sparsity. Additionally, it has been demonstrated that even leveraging pre-trained LLM models does not necessarily yield superior performance on standard time series tasks (Tan et al., 2024), highlighting the need for a sophisticated design in case of using similar architecture for such tasks. Moreover, trajectory prediction requires not only forecasting future positions but also ensuring spatial continuity and adherence to real-world constraints, a challenge that standard forecasting models do not inherently address. Our method, TrajLearn, overcomes these limitations by incorporating a constrained beam search mechanism that enforces connectivity between adjacent hexagonal blocks, thereby capturing higher-order mobility flows and complex transition patterns that reflect regional interdependencies. In this way, while transformer-based and other conventional forecasting approaches excel at modeling temporal dynamics, they fall short in handling the discrete spatial transitions and geometric constraints essential for accurate trajectory prediction.
@@ 8. Ethical Implications
Privacy and Data Protection. Using trajectory data raises concerns about individual privacy and the potential for re-identification of individuals. To protect the privacy of individuals, all datasets used in the experimental evaluation have been anonymized and aggregated. The original datasets are publicly available and are free of use (for research intent purposes), as outlined by their respective curators, and were deemed to be collected with proper informed consent and ethical approvals. Moreover, they have undergone the necessary preprocessing to meet specific research requirements. We also followed all terms and conditions of use as specified by the dataset providers, including properly attributing their work.
Other Ethical Considerations. Deep generative models pose challenges, such as the generation of plausible but incorrect data in certain scenarios. In designing and developing our model, we have made a conscientious effort to proactively address potential ethical considerations. We have been diligent in adopting responsible practices to the best of our knowledge and capacity. By doing so, we aim to ensure that deploying deep generative models for trajectory prediction aligns with societal values and prioritizes the well-being of individuals and communities.
Fairness and Biases in Trajectory Datasets. We recognize that existing trajectory datasets often contain inherent biases and may not comprehensively represent all demographic or geographic segments, and such biases can potentially lead to unequal performance across different groups. While addressing these issues is critical for developing equitable and reliable trajectory prediction systems, a detailed investigation into dataset biases is beyond the scope of this paper. We encourage future research to perform comprehensive bias assessments on publicly available trajectory datasets.
@@ 9. Conclusions
We focused on the problem of trajectory prediction and proposed TrajLearn, a trajectory deep generative model that has shown remarkable results in predicting the future path of a trajectory across various real-world datasets. Our model was trained from scratch on large-scale, higher-order mobility flow datasets that represent trajectories as sequences of hexagons on a hex-tessellated map. By utilizing higher-order mobility flows, notable data simplification while preserving essential spatial and temporal information and incorporating a constrained beam search strategy, TrajLearn achieves significant advancements over state-of-the-art methods, with performance improvements of up to 40%. Additionally, our development of mixed-resolution maps demonstrates the model’s adaptability and practicality for diverse applications, from urban planning to autonomous navigation. Our extensive empirical evaluations across three real-world datasets, Ho-Porto, Ho-Rome, and Ho-GeoLife, demonstrate that TrajLearn consistently outperforms strong baselines, with Accuracy and BLEU score improvements of up to 40%. These results not only validate the robustness of our model but also highlight its competitive performance across different resolutions, showcasing its versatility. We believe our proposed approach and model for trajectory prediction can find many valuable applications and have a broad impact. While the focus of this paper is on the development and evaluation of our current trajectory prediction framework, several promising avenues for future research lie beyond its immediate scope. For example, incorporating richer semantic and contextual information, such as detailed road network data and environmental factors, could further enhance predictive accuracy in cases where such information is available. Additionally, investigating pre-training techniques for transfer learning across diverse tasks and trajectory datasets may offer a more general framework. We encourage future work to pursue these directions to further advance the field.
Acknowledgements.
This research was supported by the Sponsor Natural Sciences and Engineering Research Council of Canada (NSERC) through an NSERC Discovery Grant (#Grant #RGPIN-2022-04586) and the NSERC CREATE training program in Dependable Internet of Things Applications (DITA).